OpenStudy (lgbasallote):

pls show me how to do this... \(\LARGE \int \frac{\sqrt{\tan x}}{1 - \sin ^2 x} dx\)

OpenStudy (lgbasallote):

it looks easy...but i have no idea what to do @_@

OpenStudy (lgbasallote):

oh wait...i saw it lol...how foolish of me to have not seen that thanks @integralsabiti

OpenStudy (anonymous):

ok u can do it :D

OpenStudy (lgbasallote):

let u = \(\sqrt {tanx }\) u^2 = tanx 2udu = sec^2 x \(\int u(2udu)\) \(\int 2u^2 du\) \(\LARGE \frac{2u^3}{3}\) \(\LARGE \frac{2\tan x \sqrt{tanx}}{3}\)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

abselutelyy

OpenStudy (lgbasallote):

how foolish of me to not have seen that lol

OpenStudy (anonymous):

It happens sometime :D